OpenStudy (kweenlexis):

A single tree produces about 2.6 × 10² lb of oxygen each year. The Amazon rainforest has about 3.9×1011 trees. About how many pounds of oxygen are produced in the rainforest each day ? Enter your answer in the box. 2.8 × 10^? lb So pretty much whats the exponent to the answer they gave, i have no idea how to do this!

OpenStudy (danjs):

\[\large \frac{ 2.6*10^2~~~lb }{ 1~~tree*yr }*\frac{ 1~~yr }{ 365~~day }*\frac{ 3.9*10^{11}~~tree }{ 1~~forrest }\]

OpenStudy (kweenlexis):

Can you explain that please?

OpenStudy (danjs):

1 tree produces that much Oxygen per Year. Divide that by 365 and you have a Tree Oxygen per day. Multiplied by how many trees are in the forrest

OpenStudy (kweenlexis):

How you find the answer though?

OpenStudy (danjs):

\[\large =\frac{ (2.6*10^2)*(3.9*10^{11} )}{ 365 }\] pounds of oxygen for the whole forrest in 1 day

OpenStudy (kweenlexis):

but how did they get 2.8?

OpenStudy (danjs):

\[\large \frac{ 10.14*10^{13} }{ 365 }=0.0278 * 10^{13}\]

OpenStudy (kweenlexis):

wouldn't it be 278? im lost

OpenStudy (danjs):

then put it in standard scientific notation, move decimal right 2 places 2.78 * 10^15

OpenStudy (kweenlexis):

how did they get 2.8?

OpenStudy (danjs):

they just rounded it 2.78 is 2.8